The level of water in a tank is 5m high. A hole of area 1cm2 is made at the bottom of the tank . The rate of leakage of water from the level is 
(g=10 m/s2).

Grade:12th pass

3 Answers

Arun Kumar IIT Delhi
9 years ago
Speed of Efflux.

But here
Since the diameter of tank isn't given.


Arun Kumar
IIT Delhi
Askiitians Faculty
ABHISHEK VASHIST
9 years ago
sir what is  A and A1
surudhi
7 years ago
The hole made at the bottom of a tank resembles an orifice and according to Toricelli’s Theorem,
The velocity of an efflux = (2gh)½ = (2x10x5)1/2 = 10 m/s.
Rate of leakage = Velocity x Area = 10 x 10-4 = 10-3 m3/s.
(Note that the area has been converted from cm2 to m2 to provide the answer in SI units)